The divergence of the vector field F = (rz, yz, z2) at the point (1,2,3) is Select one: a. O b. 9 c. 3,3,6 d. 12
Answer #1

image.png

answered by: RAMESH RAJAK
Answer #2

A F =(22,93,729 divFs Bio (22) + 3, (92) +32(22) = 7+ 2+ 27 = 42 dir F (1,2,3) = 4(3) = 1127